OpenStudy (anonymous):

Find the range and the first and third quartiles for the data set. 22, 87, 53, 43, 94, 14, 27 a. Range = 80 First quartile = 87 Third quartile = 22 b. Range = 80 First quartile = 22 Third quartile = 87 c. Range = 43 First quartile = 22 Third quartile = 87 d. Range = 94 First quartile = 43 Third quartile = 87

OpenStudy (anonymous):

4) 22, 87, 53, 43, 94, 14, 27 Sorted 14, 22, 27, 43, 53, 87, 94 Range = 94 - 14 = 80 First quartile = (22 + 27)/2 = 24.5 Third quartile = (53 + 87)/2 = 70

OpenStudy (anonymous):

That's not an option

OpenStudy (anonymous):

I would say B.

OpenStudy (anonymous):

Why B?

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

a or b cause range is 80.. id go with b tho

OpenStudy (anonymous):

I did some rounding idk why tho but I'd say B

OpenStudy (anonymous):

its b, cause  and you find the median of both.. which is ur Q's